# Causal Inference for a Hidden Treatment

### v2

@@ Abstract
In many empirical settings, directly observing a treatment variable may be infeasible although an error-prone surrogate measurement of the latter will often be available. Causal inference based solely on the observed surrogate measurement of the hidden treatment may be particularly challenging without an additional assumption or auxiliary data. To address this issue, we propose a method that carefully incorporates the surrogate measurement together with a proxy of the hidden treatment to identify its causal effect on any scale for which identification would in principle be feasible had contrary to fact the treatment been observed error-free. Beyond identification, we provide general semiparametric theory for causal effects identified using our approach, and we derive a large class of semiparametric estimators with an appealing multiple robustness property. Furthermore, we characterize the semiparametric efficiency bound for causal effects under the proposed hidden treatment model and provide a generic construction of estimators that attain this bound under standard regularity conditions. A significant obstacle to our approach is the estimation of nuisance functions involving the hidden treatment, which prevents the direct application of standard machine learning algorithms. To resolve this, we introduce a novel semiparametric EM algorithm, thus adding a practical dimension to our theoretical contributions. This methodology can be adapted to analyze a large class of causal parameters in the proposed hidden treatment model, including the population average treatment effect, the effect of treatment on the treated, quantile treatment effects, and causal effects under marginal structural models. We examine the finite-sample performance of our method using simulations and an application which aims to estimate the causal effect of Alzheimer’s disease on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative.
aDepartment of Biostatistics, Epidemiology and Informatics, University of Pennsylvania
bDepartment of Statistics and Data Science, University of Pennsylvania
Keywords: Causal inference, EM algorithm, Misclassification error, Semiparametric theory
@@ 1 Introduction
In many practical settings where one may wish to evaluate the causal effect of a hypothetical intervention from observational data, it may be challenging or impossible to directly observe the treatment of interest without error. This may arise either because of mis-classification error in ascertaining the treatment status for study units, or because the treatment is inherently latent and cannot conceivably be directly measured. Examples of the first kind abound in the health and social sciences; for instance, self-reported dietary intake in nutrition epidemiology is notoriously error-prone, making it challenging to quantify say the extent to which a person meets or exceeds the recommended daily dietary intake of a given nutrient, a potential intervention of interest. An example of the second kind arises in evaluating the potential causal impact of early onset dementia on say subsequent cognitive decline, as no gold standard currently exists for reaching a definitive diagnosis of early onset dementia (Porsteinsson et al., 2021). More broadly, unvalidated self-reported treatment may not reflect true treatment status. Hu and Lewbel (2012) illustrate this by highlighting potential discrepancies between an individual’s actual college attendance and their claimed education level. Likewise, data collected during the course of clinical practice may be incomplete for patients seeking supplemental care at an external healthcare location not captured in their primary care physician’s database. For instance, medical records for a given primary healthcare facility will typically only track on-site vaccinations, thus potentially overlooking vaccinations administered at external locations such as at patients’ local pharmacies. Therefore, a patient classified as unvaccinated in their physician’s database, may have in fact been vaccinated at their local pharmacy unbeknownst to their physician. Baum et al. (2021) illustrate the challenge of drawing valid inferences about vaccine effectiveness when vaccination status may be misclassified as described above. In all of these applications, neglecting treatment misclassification by naively evaluating the effect of error prone measurement of treatment may be subject to bias and thus may lead to invalid causal inferences.
The study of measurement error models is longstanding, driven by its broad relevance. A significant portion of this research has focused on continuous treatments, where classical measurement error, i.e., zero-mean additive and independent measurement error is routinely assumed. Various strategies have been developed to address measurement error, including distribution-based methods (e.g., deconvolution (Carroll and Hall, 1988; Fan and Truong, 1993), validation data (Sepanski and Carroll, 1993; Hu and Ridder, 2012)), auxiliary variables (e.g., repeated measurements (Delaigle et al., 2008; Li, 2002), instrumental variables (Wang and Hsiao, 2011; Schennach, 2007; Hu, 2008)), or using higher-order information (Schennach and Hu, 2013; Reiersøl, 1950). Another line of work imposes weaker assumptions, but obtaining only bounds for the parameters of interest (Klepper and Leamer, 1984; Horowitz and Manski, 1995; Schennach, 2014). Please refer to Schennach (2016) for a comprehensive review of the measurement error literature.
@@
Compared to the well-explored domain of mismeasured continuous treatments, the study of misclassified binary treatments within a causal context has been far less extensive (Bollinger, 1996). This may be partially due to the inherent nonclassical nature of misclassification error of a binary treatment, given that such error cannot generally simultaneously be additive, independent of the hidden treatment value, and mean zero, other than perhaps in certain degenerate cases (see §LABEL:supp-sec:classical_error of the Supplementary Materials for further detail). Mahajan (2006) considers the identification and estimation of nonparametric and semiparametric conditional mean regression models with a misclassified binary regressor, and leverages an instrumental variable to identify and estimate the model. Lewbel (2007) likewise uses a three-valued intrumental variable to identify and estimate the conditional average treatment effect (CATE) of a misclassified binary treatment. Hu (2008) extends the framework to a general categorical treatment subject to misclassification via a so-called matrix diagonalization technique to solve a nonlinear system of equations. Allman et al. (2009) propose a comprehensive solution for identifying the distribution of an unobserved discrete variable through multiple measurements, based on the array decomposition result in Kruskal (1977). Besides auxiliary variables, validation data is another commonly used tool, where both error-prone and error-free variables are observed on a subsample (Chen et al., 2005; Sepanski and Carroll, 1993).
@@
In the absence of auxiliary data, Chen et al. (2009) establish point identification for a nonparametric location-shift model using characteristic functions. They propose sieve maximum likelihood estimators (MLE) for certain smooth functionals and provide corresponding semiparametric efficiency theory (Shen, 1997). Compared to our work, their model cannot accommodate discrete outcomes due to the location-shift model structure, however, it obviates the need for anotherproxy or instrumental variable beyond the error prone measurement. Furthermore they provide a variational characterization of the efficiency bound for their class of functionals which may be challenging to use in practice for uncertainty quantification due to the lack of a closed-form expression. Another research strand focuses on partial identification rather than point identification. Imai and Yamamoto (2010) seek to determine bounds for the average treatment effect (ATE) when a binary treatment is measured with differential error, which allows the error to directly depend on the outcome. DiTraglia and Garcia-Jimeno (2019) extends the model in Mahajan (2006) to include an endogenous treatment, deriving both partial and point identification results under various assumptions.
@@
In this paper, we propose a general approach for the identification and semiparametric estimation of a large class of causal effects of a binary treatment subject to misclassification error, in settings where no validation data might be available, however, in addition to the error-prone treatment measurement, which throughout we refer to as surrogate treatment, one has access to a proxy variable that provides an additional exogenous source of variation associated with the hidden treatment. The main distinction we draw between a surrogate and a proxy is tied to their relative level of relevance for the hidden treatment, in the sense that the degree of misclassification error in the surrogate is known to admit a bound that ensures it outperforms a random guess (although the nature of the bound may be unknown). In contrast, misclassification error in the proxy need not be similarly bounded. Valid surrogate measurements abound in practice, often in the form of self-report treatment as in the nutrition epidemiology example given above; while proxies are also routinely available, such as in the dementia example also reported above where a specific instrument used to measure cognitive function might serve as a valid proxy of dementia onset, provided the two have an association that is not null, arguably a much lower bar than for a surrogate.
@@
Throughout the paper, we use the ATE as a running example of our proposed approach; while other causal parameters and related theoretical extensions are briefly discussed in the main body of the paper, a detailed discussion is relegated to the Supplementary Materials. Our method for identifying ATE draws on the matrix diagonalization approach detailed by Hu (2008), which we apply to a broad class of causal parameters. While the identification strategy has a long-lasting precedent, the estimation methods proposed by Mahajan (2006), Lewbel (2007), and Hu (2008) fall short of expectations. Lewbel (2007) adopts a Generalized Method of Moments (GMM) estimator for discrete covariates and employs a local-GMM approach for continuous covariates at fixed points, rendering the method less practically useful in settings with moderate to large number of continuous covariates due to the curse of dimensionality. The plug-in estimators proposed by Mahajan (2006) and Hu (2008) both follow directly from the identification procedure, making them particularly vulnerable to bias in the estimated law for the observed data distribution. This sensitivity contributes to their instability, as shown in cases where the estimated probabilities involving the hidden treatment may not always be real numbers.
@@
In contrast to the estimators mentioned above, we adopt a modern semiparametric theory framework (Bickel et al., 1993; Newey, 1990; van der Vaart, 1991), and propose a general method for constructing regular and asymptotically linear (RAL) estimators of various causal parameters of interest. Our results rely upon a novel characterization of a formal mapping between the Hilbert space of all influence functions for the causal functional of interest in the oracle model in which, contrary to fact, misclassification error is absent, to the corresponding space for all influence functions in the observed data model in which the treatment is hidden. In particular, this characterization delivers a large class of new multiply robust semiparametric estimators for a rich class of causal parameters in the hidden treatment model. Multiple robustness implies that our estimators remain consistent even if some but not all required nuisance functions are estimated consistently. An immediate implication of multiple robustness is that our estimators admit a mixed-bias property (Robins et al., 2008; Rotnitzky et al., 2021; Ghassami et al., 2022) which guarantees root-n consistency and asymptotic normality even if a subset of the nuisance functions are estimated nonparametrically at relatively slow rates, provided that other nuisance functions can be estimated at sufficiently fast rates. A unique challenge we face for inference is that nuisance functions involve the hidden treatment and therefore must be estimated with extra care, which we achieve by adopting a novel nonparametric implementation of the EM algorithm inspired by Benaglia et al. (2009). The approach appropriately targets the nuisance parameters without compromising the local robustness property of our treatment effect estimator. Importantly, our approach sidesteps the requirement of a consistent estimator of the entire observed data law and the inverse-problem challenges faced by Mahajan (2006) and Hu (2008), thereby ensuring the possibility for multiple robustness which they lack. As we establish, our estimator is provably semiparametric efficient in the important case of a binary outcome. In the more general case of a categorical outcome, we derive a large class of semiparametric estimators and characterize the efficiency bound in the hidden treatment model. Furthermore, we demonstrate that the proposed approach extends to a rich class of causal functionals for which we derive corresponding multiply robust estimators irrespective of the outcome type (i.e., whether discrete or continuous), including an approximate semiparametric efficient estimator in the continuous case.
@@
  1. 1.
    We provide a unification of previously proposed point identification techniques of a large class of causal parameters in a hidden treatment model in which a pair of proxy and surrogate measurements are available.
  2. 2.
    We present general semiparametric theory for several common causal functionals under the hidden treatment model, which entails developing a formal mapping from the oracle tangent space in which the hidden treatment is observed, to the observed data tangent space; this mapping enables a characterization of the set of influence functions for the causal parameters of interest.
  3. 3.
    We characterize the semiparametric efficiency bound for the hidden treatment model based on a closed-form expression for the efficient influence function for any causal effect of interest.
  4. 4.
    We propose a class of multiply robust estimators for functionals in the hidden treatment model which require a novel semiparametric EM algorithm for estimating nuisance functions that involve the hidden treatment. Our estimators are discussed in detail in the special case of the ATE, and evaluated via extensive empirical experiments and an application.
The rest of the paper is organized as follows. In §2, we review some common causal parameters of interest. The following sections focus primarily on the ATE as a running example to illustrate our methods. In §3, we introduce the hidden treatment model, and present the key identification results. We proceed to develop the semiparametric theory in §4, and detail the construction of a semiparametric estimator for which we provide formal asymptotic theory in §5. We illustrate the effectiveness of the method via a simulation study in §7 and a real data application on the impact of Alzheimer’s disease (AD) on hippocampal volume with imperfect diagnosis in §8. The paper concludes with a discussion in §9.
2 Examples of causal parameters and their influence functions
Let $A^{*}$ denote the hidden binary treatment, $Y$ be a binary, count, or continuous outcome, and $X$ be a vector of observed covariates. The potential outcome $Y^{a^{*}}$ represents a unit’s outcome had, possibly contrary to fact, the treatment $A^{*}$ been set to $a^{*}$. Depending on the nature of the outcome and of the scientific inquiry, a variety of causal effects might be of interest. Below we briefly review several canonical examples of causal parameters of broad interest. Throughout, we make standard positivity, consistency and no unmeasured confounding assumptions about the hidden treatment as formally stated below, allowing the exposition to focus solely on the identification and estimation challenges related to treatment misclassification. For each parameter, we briefly review their identifying functional and corresponding class of influence functions an oracle with access to $A^{*}$ might use as basis for inference.
@@ 2 Examples of causal parameters and their influence functions
Average treatment effect (ATE). The ATE measures the average additive effect of a treatment on the outcome for the entire population, $E(Y^{a^{*}=1}-Y^{a^{*}=0})$, it therefore suffices to identify $E(Y^{a^{*}=1}),a^{*}=0,1$. The ATE would in principle be identified by the oracle using the standard g-formula
@@
under the following standard causal assumptions (Robins, 1986; Hernan and Robins, 2023).
@@
i. (Ignorability) $Y^{a^{*}}\!\perp\!\!\!\perp A^{*}\mid X$ for $a^{*}=0,1$.
ii. (Positivity) $10<\textrm{pr}(A^{*}=1\mid X)<$ 1 almost surely.
iii. (Consistency) $Y^{A^{*}}=Y$ almost surely.
@@
where $f(A^{*}=1\mid X)$ is the propensity score, and $E(Y\mid A^{*},X)$ is the outcome regression function (Robins et al., 1994). The upperscript $F$ denotes the full data model where $A^{*}$ is observed. ∎
@@
The Effect of Treatment on the Treated (ETT). The ETT is defined as $\psi_{ett}=E\{Y^{a^{*}=1}-Y^{a^{*}=0}\mid A^{*}=1\}$. It is often of interest when evaluating the effect of treatments that require voluntary participation from members of the population. Under Assumption 1, the ETT is identified by $\psi_{ett}=E\{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)\mid A^{*}=1\}$. The EIF for $\psi_{ett}$ as derived in (Hahn, 1998) is given by
$\begin{split}\phi^{F}_{ett}&=\frac{\mathbb{I}(A^{*}=1)}{\textrm{pr}(A^{*}=1)}\{Y-E(Y\mid A^{*}=1,X)\}\\ &\quad-\frac{\mathbb{I}(A^{*}=0)}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}\{Y-E(Y\mid A^{*}=0,X)\}\\ &\quad+\frac{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)-\psi_{ett}}{\textrm{pr}(A^{*}=1)}\textrm{pr}(A^{*}=1\mid X).\end{split}$ (3)
@@
Quantile treatment effect (QTE). Quantile treatment effects examine the impact of a treatment on a given quantile of the outcome (e.g., the median), rather than focusing on the treatment effect on the average outcome. Suppose we are interested in the $\gamma$-quantile of a continuous potential outcome $Y^{a^{*}}$, i.e., the $\theta_{1}^{a^{*}}$ such that $\textrm{pr}(Y^{a^{*}}\leq\theta_{1}^{a^{*}})=\gamma$. Under Assumption 1, the QTE can be identified. The efficient influence function for $\theta_{1}$ is proportional to
@@
where $\eta_{1}(X;\theta_{1}^{a^{*}})=\textrm{pr}(Y\leq\theta_{1}^{a^{*}}\mid X,A^{*}=a^{*})$ and $\eta_{2}(a^{*};X)=\textrm{pr}(A^{*}=a^{*}\mid X)$. ∎
@@
where $h(X)$ is any user-specified function with finite variance. ∎
@@
$\begin{split}\phi^{F}_{msm}&=Wh(A^{*},V)\left\{Y-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}\\ &\quad-Wh(A^{*},V)\left\{E(Y\mid A^{*},X)-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}\\ &\quad+\sum_{a^{*}}(-1)^{1-a^{*}}h(a^{*},V)\left\{E(Y\mid A^{*}=a^{*},X)-g^{-1}(\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\},\end{split}$ (6)
where $W=1/\textrm{pr}(A^{*}\mid X)$, and $h(A^{*},V)$ is a user-specified function with the same dimension as $(\beta_{0},\beta_{1},\beta_{2}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. ∎
Because influence functions have zero mean only at the truth, in principle they provide a convenient approach for constructing semiparametric estimators. For instance, in Example 1, by solving the estimating equation $\mathbb{P}_{n}\{\widehat{\phi}_{a^{*}}^{F}(\widehat{\psi}_{a^{*}})\}=0$ based on an estimate of the efficient influence function (2), one recovers the well-known augmented-inverse-probability-weighted estimator (AIPW):
@@
where $\hat{f}(A^{\ast}=a^{*}\mid X)$ and $\hat{E}(Y\mid a^{*},X)$ are estimated propensity score and outcome regression function, respectively. This estimator is known to be doubly robust (Scharfstein et al., 1999), in that $\widehat{\psi}_{a^{*}}$ is consistent if either $\hat{f}(A^{\ast}=a^{*}\mid X)$ or $\hat{E}(Y\mid a^{*},X)$ is consistent, but not necessarily both. When both are consistent, the estimator achieves the semiparametric efficiency bound for the nonparametric model.
As $A^{*}$ is not directly observed in our setting, the estimator (7) is of course not feasible. In fact, $\psi_{a^{*}}$ cannot be identified without an additional condition. In the next section, we provide a general approach towards identification correcting for the misclassification error by leveraging two types of auxiliary variables: a surrogate $A$ and a proxy $Z$ of the hidden treatment. Instead of $A^{*}$, as will routinely be the case in practice, we presume that a misclassified measurement $A$ of the treatment variable is available; e.g., in case $A^{*}$ is whether a patient is currently taking a certain medication, then $A$ might be, either self-reported or extracted from medical records both of which might be subject to error, say if the patient is non-compliant and therefore the recorded treatment does not match that actually taken by the patient. Throughout, we refer to such an error prone measurement as a surrogate variable to the extent that it reflects a genuine attempt to measure the hidden treatment. In contrast, a proxy $Z$ of treatment must be associated with the treatment, although it may itself not be an explicit measurement of the treatment. Thus, a proxy is only required to be a correlate of the hidden treatment, while the surrogate is a misclassified treatment measurement. These definitions are formalized in §3 below.

### v3

@@ Abstract
In many empirical settings, directly observing a treatment variable may be infeasible although an error-prone surrogate measurement of the latter will often be available. Causal inference based solely on the surrogate measurement is particularly challenging without validation data. We propose a method that obviates the need for validation data by carefully incorporating the surrogate measurement with a proxy of the hidden treatment to obtain nonparametric identification of several causal effects of interest, including the population average treatment effect, the effect of treatment on the treated, quantile treatment effects, and causal effects under marginal structural models. For inference, we provide general semiparametric theory for causal effects identified using our approach and derive a large class of semiparametric efficient estimators with an appealing multiple robustness property. A significant obstacle to our approach is the estimation of nuisance functions which involve the hidden treatment therefore preventing the direct use of standard machine learning algorithms, which we resolve by introducing a novel semiparametric EM algorithm. We examine the finite-sample performance of our method using simulations and an application which aims to estimate the causal effect of Alzheimer’s disease on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative.
Keywords: EM algorithm, Misclassification error, Semiparametric theory
@@ 1 Introduction
In many practical settings where one may wish to evaluate the causal effect of a hypothetical intervention from observational data, it may be challenging or impossible to directly observe the treatment of interest without error. This may arise either because of mis-classification error in ascertaining the treatment status, or because the treatment is inherently latent and cannot conceivably be directly measured. Examples of the first kind abound in health and social sciences; for instance, self-reported dietary intake in nutrition epidemiology is notoriously error-prone, making it challenging to quantify the extent to which a person meets or exceeds recommended daily dietary intake of a given nutrient, a potential intervention of interest. An example of the second kind arises in evaluating the potential causal impact of early onset dementia on subsequent cognitive decline, as no gold standard currently exists for reaching a definitive diagnosis of early onset dementia (Porsteinsson et al.,, 2021). More broadly, unvalidated self-reported treatment may not reflect true treatment status. Hu and Lewbel, (2012) illustrate this by highlighting potential discrepancies between an individual’s actual college attendance and their claimed education level.
@@
The study of measurement error models is longstanding, driven by its broad relevance. A significant portion of this research has focused on continuous treatments, where classical measurement error is routinely assumed. Various strategies have been developed to address measurement error, and Schennach, (2016) provides a comprehensive review of existing measurement error literature. Compared to the well-explored domain of mismeasured continuous treatments, the study of misclassified binary treatments within a causal context has been far less extensive (Bollinger,, 1996). This may stem from the inherent nonclassical nature of misclassification error of a binary treatment, as such error cannot generally be additive, independent of the hidden treatment value, and mean zero simultaneously, other than perhaps in certain degenerate cases (see §LABEL:supp-sec:classical_error of the Supplementary Materials for details). Mahajan, (2006) considers the identification and estimation of nonparametric and semiparametric conditional mean regression models with a misclassified binary regressor, and leverages an instrumental variable to identify and estimate the model. Lewbel, (2007) likewise uses a three-valued intrumental variable to identify and estimate the conditional average treatment effect (CATE) of a misclassified binary treatment. Hu, (2008) extends the framework to a general categorical treatment subject to misclassification via a so-called matrix diagonalization technique to solve a nonlinear system of equations. Allman et al., (2009) propose a comprehensive solution for identifying the distribution of an unobserved discrete variable through multiple measurements, based on array decomposition results. Validation data is another commonly used tool, where both error-prone and error-free variables are observed on a subsample (Chen et al.,, 2005; Sepanski and Carroll,, 1993).
@@
In the absence of auxiliary data, Chen et al., (2009) establish point identification for a nonparametric location-shift model using characteristic functions. They propose sieve maximum likelihood estimators for certain smooth functionals and provide corresponding semiparametric efficiency theory (Shen,, 1997). They also provide a variational characterization of the efficiency bound for their functionals which may be challenging to use in practice for uncertainty quantification due to the lack of a closed-form expression. Another research strand focuses on partial identification rather than point identification. Imai and Yamamoto, (2010) seek to determine bounds for the average treatment effect (ATE) when a binary treatment is measured with differential error, which allows the error to depend on the outcome. DiTraglia and Garcia-Jimeno, (2019) extends the model in Mahajan, (2006) to accommodate endogeneity of the treatment, establishing conditions for partial identification and point identification.
@@
In this paper, we propose a general approach for the identification and semiparametric estimation of a large class of causal effects of a misclassified binary treatment, even in the absence of validation data. However, in addition to the error-prone treatment measurement, referred to as surrogate treatment, one has access to a proxy variable that provides an additional exogenous source of variation associated with the hidden treatment. The main distinction between a surrogate and a proxy lies in their relative level of relevance to the hidden treatment. The degree of misclassification error in the surrogate is known to admit a bound that ensures it outperforms a random guess (although the nature of the bound may be unknown). In contrast, misclassification error in the proxy needs not be similarly bounded. Valid surrogate measurements abound in practice, often in the form of self-report treatment as in the nutrition epidemiology example given above. Proxies are also routinely available; for example, a cognitive function test may serve as a proxy for dementia onset if the two have a non-null association, a less stringent requirement than for a surrogate.
@@
Throughout the paper, we use the ATE as a running example of our proposed approach; while other causal parameters and related theoretical extensions are briefly discussed in the main body of the paper, a detailed discussion is relegated to the Supplementary Materials. Our method for identifying ATE draws on the matrix diagonalization approach detailed by Hu, (2008), which we apply to a broad class of causal parameters. While the identification strategy has a longstanding precedent, the estimation methods proposed by Mahajan, (2006), Lewbel, (2007), and Hu, (2008) have several limitations. Mainly, Lewbel, (2007) adopts a Generalized Method of Moments (GMM) estimator for discrete covariates and employs a local-GMM approach for continuous covariates at a fixed point, limiting the practical utility of their approach. The plug-in estimators proposed by Mahajan, (2006) and Hu, (2008) both follow directly from the identification procedure, making them particularly vulnerable to bias in the estimated law for the observed data distribution. This sensitivity contributes to their instability, as shown in cases where the estimated probabilities involving the hidden treatment may not always be real numbers.
@@
In contrast to the estimators mentioned above, we adopt a modern semiparametric theory framework (Bickel et al.,, 1993; Newey,, 1990; van der Vaart,, 1991) to develop a general method for constructing regular and asymptotically linear (RAL) estimators of various causal parameters of interest. Our results rely upon a novel characterization of a formal mapping between the Hilbert space of all influence functions for the causal functional of interest in the idealized oracle model where misclassification error is absent, and the corresponding space of all influence functions in the observed data model in which the treatment is hidden. This characterization delivers a large class of new multiply robust semiparametric estimators for a rich class of causal parameters in the hidden treatment model. Multiple robustness ensures that our estimators remain consistent even if some but not all required nuisance functions are estimated consistently. This implies that our estimators admit a mixed-bias property (Robins et al.,, 2008; Rotnitzky et al.,, 2021; Ghassami et al.,, 2022) which guarantees root-n consistency and asymptotic normality even if some nuisance functions are estimated nonparametrically at relatively slow rates, provided that other nuisance functions can be estimated at sufficiently fast rates. A unique challenge we face for inference is that nuisance functions involve the hidden treatment, requiring careful estimation. We address this by adopting a novel semiparametric implementation of the EM algorithm inspired by Benaglia et al., (2009). The approach appropriately targets the nuisance parameters without compromising the local robustness property of our treatment effect estimator. Importantly, our approach avoids the need for a consistent estimator of the entire observed data law and bypasses the inverse-problem challenges faced by Mahajan, (2006) and Hu, (2008), thereby ensuring the possibility for multiple robustness which they lack. As we establish, our estimator is semiparametric efficient in the important case of a binary outcome. For more general categorical outcomes, we derive a large class of semiparametric estimators and characterize the efficiency bound in the hidden treatment model. Furthermore, we extend our approach to a rich class of causal functionals, providing multiply robust estimators for both discrete and continuous outcomes, including an approximate semiparametric efficient estimator in the continuous case.
@@
1. We unify existing point identification techniques for a large class of causal parameters in hidden treatment models where both proxy and surrogate measurements are available.
2. We present semiparametric theory for several common causal functionals under the hidden treatment model by developing a formal mapping from the oracle tangent space where the hidden treatment is observed, to the observed data tangent space. This mapping enables a characterization of the set of influence functions for the causal parameters of interest.
3. We characterize the semiparametric efficiency bound for the hidden treatment model based on a closed-form expression for the efficient influence function for any causal parameter.
4. We propose a class of multiply robust estimators for functionals in the hidden treatment model which require a novel semiparametric EM algorithm for estimating nuisance functions involving the hidden treatment. Our estimators are discussed in detail in the special case of the ATE, and evaluated via empirical experiments and an application.
The rest of the paper is organized as follows. §2 reviews some common causal parameters of interest. In §3, we introduce the hidden treatment model and present key identification results. §4 develops the semiparametric theory for the ATE, and §5 details the construction of a semiparametric estimator along with its formal asymptotic theory. §6 extends these results to other causal functionals. We illustrate the effectiveness of the method via a simulation study in §7 and an application on the impact of Alzheimer’s disease on hippocampal volume with imperfect diagnosis in §8. The paper concludes with a discussion in §9.
2 Examples of causal parameters
Let $A^{*}$ denote the hidden binary treatment, $Y$ be a binary, count, or continuous outcome, and $X$ be a vector of observed covariates. The potential outcome $Y^{a^{*}}$ represents a unit’s outcome had, possibly contrary to fact, the treatment $A^{*}$ been set to $a^{*}$. Depending on the nature of the outcome and of the scientific inquiry, a variety of causal effects might be of interest. Below we briefly review several examples of causal parameters of broad interest. Throughout, we make standard positivity, consistency and no unmeasured confounding assumptions about the hidden treatment as formally stated below, allowing the exposition to focus solely on the identification and estimation challenges related to treatment misclassification. For each parameter, we briefly review its identifying functional and corresponding class of influence functions that an oracle with access to $A^{*}$ might use as basis for inference.
@@ 2 Examples of causal parameters
Average treatment effect (ATE). The ATE measures the average additive effect of a treatment on the outcome, $E(Y^{a^{*}=1}-Y^{a^{*}=0})$. It therefore suffices to identify $E(Y^{a^{*}}),a^{*}=0,1$. The ATE can be identified by the oracle using the standard g-formula
@@
under the following standard causal assumptions (Robins,, 1986).
@@
i. (Ignorability) $Y^{a^{*}}\!\perp\!\!\!\perp A^{*}\mid X$ for $a^{*}=0,1$; ii. (Positivity) $10<\textrm{pr}(A^{*}=1\mid X)<$ 1 almost surely; iii. (Consistency) $Y^{A^{*}}=Y$ almost surely.
@@
where $f(A^{*}=a^{*}\mid X)$ is the propensity score and $E(Y\mid A^{*},X)$ is the outcome regression function (Robins et al.,, 1994). The upperscript $F$ denotes the full data model where $A^{*}$ is observed.
Remark 1.
For categorical $Y$, the parameters of interest would be $E\{\mathbb{I}(Y^{a^{*}}=y)\}$ where $y$ is any category $Y$ can assume. Its EIF can be obtained by replacing $Y$ with $\mathbb{I}(Y=y)$ in (2).
@@
The Effect of Treatment on the Treated (ETT). The ETT is defined as $\psi_{ett}=E\{Y^{a^{*}=1}-Y^{a^{*}=0}\mid A^{*}=1\}$. It is often of interest when evaluating the effect of treatments that require voluntary participation from members of the population. Under Assumption 1, the ETT is identified by $\psi_{ett}=E\{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)\mid A^{*}=1\}$. The EIF for $\psi_{ett}$ as derived in Hahn, (1998) is given by
$\phi^{F}_{ett}$ $=\frac{\mathbb{I}(A^{*}=1)}{\textrm{pr}(A^{*}=1)}\{Y-E(Y\mid A^{*}=1,X)\}-\frac{\mathbb{I}(A^{*}=0)}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}\{Y-E(Y\mid A^{*}=0,X)\}$
$\quad+\frac{E(Y\mid A^{*}=1,X)-E(Y\mid A^{*}=0,X)-\psi_{ett}}{\textrm{pr}(A^{*}=1)}\textrm{pr}(A^{*}=1\mid X).$ (3)
@@
Quantile treatment effect (QTE). Quantile treatment effects examine the impact of a treatment on a given quantile of the outcome (e.g., the median), rather than focusing on the treatment effect on the average outcome. Suppose we are interested in the $\gamma$-quantile of a continuous potential outcome $Y^{a^{*}}$, i.e., the $\theta_{1}^{a^{*}}$ such that $\textrm{pr}(Y^{a^{*}}\leq\theta_{1}^{a^{*}})=\gamma$. Under Assumption 1, the QTE can be identified. The EIF for $\theta_{1}$ is proportional to
@@
where $\eta_{1}(X;\theta_{1}^{a^{*}})=\textrm{pr}(Y\leq\theta_{1}^{a^{*}}\mid X,A^{*}=a^{*})$ and $\eta_{2}(a^{*};X)=\textrm{pr}(A^{*}=a^{*}\mid X)$.
@@
where $h(X)$ is any user-specified function with finite variance.
@@
$\phi^{F}_{msm}$ $=Wh(A^{*},V)\left\{Y-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}-Wh(A^{*},V)\left\{E(Y\mid A^{*},X)-g^{-1}(\beta_{0}+\beta_{1}A^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\}$
$\quad+\sum_{a^{*}}(-1)^{1-a^{*}}h(a^{*},V)\left\{E(Y\mid A^{*}=a^{*},X)-g^{-1}(\beta_{0}+\beta_{1}a^{*}+\beta_{2}^{\mathrm{\scriptscriptstyle T}}V)\right\},$ (6)
where $W=1/\textrm{pr}(A^{*}\mid X)$, and $h(A^{*},V)$ is a user-specified function with the same dimension as $(\beta_{0},\beta_{1},\beta_{2}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
Because influence functions have zero mean only at the truth, in principle they provide a convenient approach for constructing semiparametric estimators. For instance, in Example 1, by solving the estimating equation $\mathbb{P}_{n}\{\widehat{\phi}_{a^{*}}^{F}(\widehat{\psi}_{a^{*}})\}=0$ based on an estimate of the EIF (2), one recovers the well-known augmented-inverse-probability-weighted estimator (AIPW):
@@
where $\hat{f}(A^{\ast}=a^{*}\mid X)$ and $\hat{E}(Y\mid a^{*},X)$ are estimated propensity score and outcome regression function, respectively. This estimator is known to be doubly robust (Scharfstein et al.,, 1999), in that $\widehat{\psi}_{a^{*}}$ is consistent if either $\hat{f}(A^{\ast}=a^{*}\mid X)$ or $\hat{E}(Y\mid a^{*},X)$ is consistent, but not necessarily both. When both are consistent, the estimator achieves the semiparametric efficiency bound for the nonparametric model.
As $A^{*}$ is not directly observed in our setting, the estimator (7) is not feasible, and $\psi_{a^{*}}$ cannot be identified without additional conditions. In the next section, we provide a general approach towards identification correcting for the misclassification error using two auxiliary variables: a surrogate $A$ and a proxy $Z$ of the hidden treatment. Instead of observing $A^{*}$, we assume access to a misclassified measurement $A$. For example, if $A^{*}$ indicates whether a patient is taking a medication, $A$ could be self-reported or extracted from medical records, both prone to errors, such as when a non-compliant patient’s recorded treatment does not match their actual intake. We refer to such an error-prone measurement as a surrogate variable as it attempts to measure the hidden treatment. In contrast, a proxy $Z$ of treatment is only required to be associated with the hidden treatment, although it may not be an explicit measurement of the treatment. These definitions are formalized in §3.